OpenStudy (anonymous):

Converting 98.1827 degrees to radians in fractions? (example 32.73 degrees = 2(pi)/11

OpenStudy (anonymous):

\[Angle(radians)=Angle(degrees)\times\frac{2\pi}{360}\]